OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14246 — TPIX
Incident Report for Network & Infrastructure
Resolved
We detected some BGP flaps during the night and this morning at the TPIX.

We contacted them to have more information about the issue.

This task will be updated as soon as we know more.

Update(s):

Date: 2015-08-05 13:41:02 UTC
We are redirecting the traffic to TPIX as IX confirmed that the issue has been resolved.

Date: 2015-08-04 10:49:37 UTC
We've redirected the traffic to and from TPIX while we wait for IX to confirm that everything is fixed.

Date: 2015-08-04 10:48:58 UTC
This is a strange issue. There was a loss of light at
about 1.30am CEST. During the night the traffic decreased
between 3.40am - 4.10am and 6.20-7.00am CEST. At about
9.30 we lost a BGP session (as hold-time
expired) and after this the traffic went back to normal.
Posted Aug 04, 2015 - 09:16 UTC